OpenStudy (anonymous):

Use implicit differentiation to find the slope of the tangent line to the curve {y/{x - 7 y} = x^{9} + 3 at the point ( 1, (4/29) ).

OpenStudy (anonymous):

\[\frac{y}{x-7y}=x^9+3\]?

OpenStudy (anonymous):

yes that is the curve

OpenStudy (anonymous):

take the derivative implicitly. you need the quotient rule on the left. you should get \[\frac{(x-7y)y'-y(1-7y')}{(x-7y)^2}=9x^8\]

OpenStudy (amistre64):

id kill the fraction and product rule it for a cleaner feel and a fresher scent

OpenStudy (anonymous):

do some algebra, plug in the numbers and solve for y'

OpenStudy (anonymous):

better idea!

OpenStudy (anonymous):

from grease to shine in half the time...

OpenStudy (amistre64):

\[\frac{y}{x-7y}=x^9+3\] \[y=(x^9+3)(x-7y)\] \[y'=(x^9+3)'(x-7y)+(x^9+3)(x-7y)'\] \[y'=(9x^8)(x-7y)+(x^9+3)(x'-7y')\]

OpenStudy (anonymous):

i definitely liked your second method better lol thank you!

OpenStudy (anonymous):

Thank you for explaining the first method too, figuring how to do it that way was killing me until you explained it

OpenStudy (amistre64):

the quotient rule is inherently a nightmare ...